OpenStudy (anonymous):

I REALLY NEED HELP WITH THIS QUESTION , SOMEBODY PLEASE HELP ME , MEDAL AND FAN Graph an ellipse with foci at (0, ±2square root of 10 ) and x-intercept at (9, 0) using graphing technology.

OpenStudy (rock_mit182):

a^

OpenStudy (rock_mit182):

case A :\[\frac{(x-h)^{2} }{ b } + \frac{ (y-k)^{2} }{ a } = 1\] Ellipse's equation